Three coins are tossed. Let T be the number of tails that occur.

a. Construct the probability distribution for the random variable T.

b. Construct the probability histogram


1
Expert's answer
2021-03-30T13:24:38-0400

а. We will assume that the probability of getting heads and tails is the same:

"p = q = \frac{1}{2}"

Using Bernoulli's formula, we find the probabilities that 0, 1, 2, and 3 tails will appear:

"p(0) = {q^3} = {\left( {\frac{1}{2}} \right)^3} = \frac{1}{8}"

"p(1) = C_3^1p{q^2} = 3 \cdot {\left( {\frac{1}{2}} \right)^3} = \frac{3}{8}"

"p(2) = C_3^2{p^2}q = 3 \cdot {\left( {\frac{1}{2}} \right)^3} = \frac{3}{8}"

"p(3) = {p^3} = {\left( {\frac{1}{2}} \right)^3} = \frac{1}{8}"

We get the probability distribution:

"\begin{matrix}
T&0&1&2&3\\
p&{\frac{1}{8}}&{\frac{3}{8}}&{\frac{3}{8}}&{\frac{1}{8}}
\end{matrix}"

b. Construct the probability histogram: